OpenStudy (anonymous):

Which of the following is a form of the law of cosines? A) a^2=b^2+c^2+2bc*cos A B) a^2=b^2-c^2+2bc*cos A C) b=a^2+c^2-2ac*cos B D) b^2=a^2+c^2-2ac*cos B

OpenStudy (usukidoll):

Law of cosines looks like this \[cosA = \frac{b^2+c^2-a^2}{2bc}\] \[cosB = \frac{a^2+c^2-b^2}{2ac}\] \[cosC=\frac{a^2+b^2-c^2}{2ab}\] to have \[a^2 \] or \[b^2\] by itself we have to manipulate this a bit.

OpenStudy (usukidoll):

forgot the \[c^2 \] case , but based on the choices you're given, I don't think you need it.

OpenStudy (anonymous):

is it answer A?

OpenStudy (anonymous):

or C